A proton in a linear accelerator has a de Broglie wavelength of (1.43x102) pm. What is the speed of the proton (in m/s)?

1
Expert's answer
2021-10-07T03:10:01-0400

From de Brogile wavelength, "\lambda=\frac{h}{mv}"

"velocity, v=h/\lambda m"

mass of proton, "m=1.673\times10^{-27}kg"

Planck's constant, "h=6.626\times 10^{-34}m^2kg/s"

"wavelength, \lambda=1.143\times 10^{-7}m"

velocity, "v=\frac{6.626\times 10^{-34}}{1.43\times 10^{-7} \times 1.673\times 10^{-27}}=2.7696m/s"